# Variational Calculation of the Hyperfine Stark Effect in Atomic 87Rb, 133Cs, and 169Tm

### v1

@@ III.4.1 ${}^{2}F_{7/2}(F=4)$${}^{2}F_{5/2}(F=3)$ clock transition
The final value for $\Delta\alpha$ is calculated by using the individual result from model QZ/SD41_10au with the largest number of correlated electrons as a base value and adding a correction due to taking into account full triple excitations among the $15$ outermost electrons, formally,
[…]
Interestingly, the present final result for $\Delta\alpha$ differs from the one measured and calculated in Ref. NatComm_Tm_2019 , even if assigned uncertainties are taken into account. Since the differential polarizability amounts to only about $0.2$% of an individual level polarizability, the former is by its nature very difficult to calculate accurately. The cited result for $\Delta\alpha$ from Ref. NatComm_Tm_2019 is also given with a relatively large uncertainty, in part due to uncertainties in employed transition probabilities that have been estimated ibidem.
Nevertheless, the present series of calculations allows for some conclusions on this problem. The major correction to the base value comes from higher excitation ranks in the wavefunction expansion (triples terms) and is negative. The importance of this correction is not surprising, as these higher excitations can be interpreted as the mixing in of excited states fulfilling $\Delta J=0$ into the atomic wavefunctions with defined $J$, thus leading to a differential effect for the $\alpha_{J}$ results. It is, therefore, to be expected that including higher excitation ranks (quadruples etc.) will lead to a more negative value for $\Delta\alpha$ and thus to an even greater discrepancy with the result by Golovizin et al.
There is good correspondence of the present most accurate result with high-level literature results for $\alpha_{J}$, but this agreement does not yield a constraint on the uncertainty of the much smaller differential polarizability $\Delta\alpha$.
III.4.2 ${}^{2}F_{7/2}(F=4)$${}^{2}F_{7/2}(F=3)$ transition
When the hyperfine transition among the two microstates belonging to the same ${}^{2}F_{7/2}$ ground level is considered, Eq. (7) is relevant. Results for the ground-term hyperfine Stark coefficient are compiled in Table 4.
Table 4: Stark hyperfine frequency shift $k$ calculated according to Eq. (5) for 169Tm ($4f_{7/2}$), $I=0.5$, $\mu=-0.2310(15)\mu_{N},E_{z}=0.001$ a.u., hyperfine level quantum numbers for the ground-state transition are $(J=7/2)F_{l}=4$, $(J=7/2)F_{u}=3$. All results are given for the electronic microstates with $M_{J}=1/2$ except where stated otherwise.
Basis set/Model $k$ [$10^{-14}$ Hz/((V/m)2)]
TZ/DCHF $-1.38$
TZ/SD15_6au $-6.36$
TZ/SD15_10au $-6.85$
TZ/SD15_30au $-7.80$
TZ/SD15_60au $-7.69$
QZ/DCHF($M_{J}=1/2$) $-1.41$
QZ/DCHF($M_{J}=3/2$) $-1.20$
QZ/DCHF($M_{J}=5/2$) $-0.57$
QZ/SD15_6au $-5.22$
QZ/SD15_10au $-7.32$
QZ/SD15_30au $-7.76$
QZ/SD15_60au $-7.67$
QZ/SDT15_10au $2.08$
QZ/S8_SD23_10au $4.93$
QZ/SD23_10au $4.98$
QZ/SD23_30au $4.28$
QZ/SD33_10au $4.47$
QZ/SD41_10au $3.94$
QZ+/DCHF $-1.41$
QZ+/SD15_5au $-5.12$
QZ+/SD15_10au $-7.00$
k(final) ${\boldsymbol{13.0\pm 10.0}}$
Concerning atomic basis sets, DCHF results are seen to be converged already for the QZ set. When correlation effects among the $15$ outermost electrons are included the largest densified basis set QZ+ gives a correction of $4-5$% to the hyperfine Stark coefficient which will be considered in the final evaluation. Additional correlation effects, on the other hand, are investigated using the computationally less expensive QZ set.
DCHF results for the $M_{J}=1/2$ microstates show the largest hyperfine Stark effect, whereas the effect of the electric field reduces as $M_{J}$ is increased (QZ/DCHF models). This can be explained as follows. The respective (leading, if a model including electron correlation effects is considered) electronic Slater determinants can be written in terms of occupied spinors in the $4f$ shell as
$\left|J=\frac{7}{2},M_{J}=\frac{1}{2}\right>$ $=$ $\left|4f_{5/2,5/2\ldots-5/2};4f_{7/2,7/2};4f_{7/2,-7/2};4f_{7/2,5/2};4f_{7/2,-5/2};4f_{7/2,3/2};4f_{7/2,-3/2};4f_{7/2,1/2}\right>$
$\left|J=\frac{7}{2},M_{J}=\frac{3}{2}\right>$ $=$ $\left|4f_{5/2,5/2\ldots-5/2};4f_{7/2,7/2};4f_{7/2,-7/2};4f_{7/2,5/2};4f_{7/2,-5/2};4f_{7/2,3/2};4f_{7/2,1/2};4f_{7/2,-1/2}\right>$
$\left|J=\frac{7}{2},M_{J}=\frac{5}{2}\right>$ $=$ $\left|4f_{5/2,5/2\ldots-5/2};4f_{7/2,7/2};4f_{7/2,-7/2};4f_{7/2,5/2};4f_{7/2,3/2};4f_{7/2,-3/2};4f_{7/2,1/2};4f_{7/2,-1/2}\right>.$
[…]
$A_{J,M_{J}}$ $=$ $4f_{j,m_{j}}\right|\hat{H}_{\text{HF}}\left|4f_{j,m_{j}}\right>$ (16)
${\tilde{A}}_{J,M_{J}}$ $=$ ${\tilde{4f}}_{j,m_{j}}\right|\hat{H}_{\text{HF}}\left|{\tilde{4f}}_{j,m_{j}}\right>$ (17)
where $\tilde{X}$ denotes a quantity perturbed by the external electric field. According to Eq. (8) the energy of the lower hyperfine level is
$\varepsilon_{F,J,M_{J}}=\varepsilon_{4,\frac{7}{2},M_{J}}\approx\frac{1}{2}\,\left[20-\frac{3}{4}-\frac{63}{4}\right]\,A_{\frac{7}{2},M_{J}}=\frac{7}{4}A_{\frac{7}{2},M_{J}}$ (18)
and the energy of the upper hyperfine level is
$\varepsilon_{3,\frac{7}{2},M_{J}}\approx\frac{1}{2}\,\left[12-\frac{3}{4}-\frac{63}{4}\right]\,A_{\frac{7}{2},M_{J}}=-\frac{9}{2}A_{\frac{7}{2},M_{J}}$ (19)
Correspondingly, when $E_{\text{ext}}$ is included,
$\tilde{\varepsilon}_{4,\frac{7}{2},M_{J}}\approx\frac{7}{4}\tilde{A}_{\frac{7}{2},M_{J}}$ (20)
$\tilde{\varepsilon}_{3,\frac{7}{2},M_{J}}\approx-\frac{9}{2}\tilde{A}_{\frac{7}{2},M_{J}}$ (21)
Note that $\varepsilon_{4,\frac{7}{2},M_{J}}<\varepsilon_{3,\frac{7}{2},M_{J}}$ since $\tilde{A}_{\frac{7}{2},M_{J}}<0$ ${}~{}\forall M_{J}$ as numerical results show which makes the level with $F=4$ the lower level.
Then the Stark coefficient for states with defined $M_{J}$ results from Eq. (5) as
$k_{M_{J}}$ $=$ $\frac{\tilde{\varepsilon}_{3,\frac{7}{2},M_{J}}-\varepsilon_{3,\frac{7}{2},M_{J}}-\tilde{\varepsilon}_{4,\frac{7}{2},M_{J}}+\varepsilon_{4,\frac{7}{2},M_{J}}}{hE_{\text{ext}}^{2}}$ (22)
$=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[-\frac{9}{2}\left(\tilde{A}_{\frac{7}{2},M_{J}}-A_{\frac{7}{2},M_{J}}\right)-\frac{7}{4}\left(\tilde{A}_{\frac{7}{2},M_{J}}-A_{\frac{7}{2},M_{J}}\right)\right]$
$=$ $\frac{1}{hE_{\text{ext}}^{2}}\left(-\frac{25}{4}\right)\left(\tilde{A}_{\frac{7}{2},M_{J}}-A_{\frac{7}{2},M_{J}}\right)$
Numerical evaluation of this expression shows that indeed $\left|k_{\frac{1}{2}}\right|>\left|k_{\frac{3}{2}}\right|>\left|k_{\frac{5}{2}}\right|$ which in essence is due to the magnitude of $\tilde{A}_{J,M_{J}}-A_{J,M_{J}}$ decreasing with increasing value of $M_{J}$. This, in turn, is readily explained by the fact that the orbital amplitude of the $4f$ wavefunction in field direction ($z$ axis) is maximal for minimal $M_{J}$, thus leading to the greatest field-dependent shift when $M_{J}=\frac{1}{2}$.
Including full triple excitations among the $15$ valence electrons (model SDT15) yields a relatively sizeable positive correction that brings the hyperfine Stark coefficient back close to the DCHF value (however, with opposite sign). It is remarkable that this correction counteracts the correlation effect from single and double excitations. Since the model SDT15 comprises nearly one billion expansion terms it is computationally impossible with the present code to investigate the effects of even higher excitation ranks than full triple replacements.
Accounting for electron correlation effects among and with the eight electrons in the Tm $5s,5p$ shells (model SD23_30au) yields a sizeable positive correction of $+155$% relative to the model SD15 that in addition changes the sign of $k$. It is not surprising that such physical effects are important since the hyperfine structure in an external $E$ field is dominated by the atomic structure in $s$ and $p$ states. In accord with this reasoning, the inclusion of correlation effects with and among the $4d$ electrons (model SD33) results in an only relatively minor correction of about $-10$%. Adding the $4s$ and $4p$ electrons (model SD41)yields a slightly more important contribution of about $-12$%, again in line with the above arguments.
The final value for the thulium hyperfine Stark shift can now be obtained from the above series of calculations. The result with $23$ electrons considered in the correlation treatment and the highest cutoff value for the virtual spinors serves as the base value. This base value is corrected for by the effects of densifying the basis set, correlating inner-shell electrons and taking triple excitations from the $15$ outermost electrons into account. Formally,
[…]
which is given in Table 4. This final result is greater than any of the individual results, both as a plain value and in magnitude, essentially because the base value is positive and the largest correction comes from higher excitation ranks which is strongly positive.
The uncertainty on this final result is also determined from the presented series of calculations. An individual uncertainty for an effect is determined as the difference between the most accurate and the second most accurate model for that effect. Then, the uncertainties are $5$% for the atomic basis set, $70$% for higher excitation ranks, and $10$% for inner-shell correlation effects. From this, a conservative estimate of ±$\pm 80$% is derived for the final value, since the considered corrections partially act with opposing signs.
@@ IV Conclusions
The presented variational relativistic configuration-interaction calculations of the hyperfine Stark coefficient are shown to be accurate in comparison with reliable literature results on the order of $<7$% for atomic 87Rb and 133Cs. The corresponding coefficient for 169Tm is significantly more difficult to calculate and the presented result is given with a consequently rather large uncertainty margin. The difficulties in obtaining accurate electron correlation effects for relevant properties of 169Tm have also been encountered in the calculation of its ground-state electric quadrupole moment Fleig_PhysRevA.107.032816 . Nevertheless, the hyperfine Stark coefficient for the ground-level hyperfine transition is smaller than $2.3\times 10^{-13}$ [Hz/((V/m)2)] – even in the $M_{J}=\frac{1}{2}$ levels where its value is greatest – which is about three orders of magnitude smaller than same coefficient for the hyperfine transitions in 87Rb and 133Cs. The present final result for the differential polarizability of the proposed clock transition in Ref. NatComm_Tm_2019 differs from the one measured and calculated in that paper by about a factor of 2. The origin of this discrepancy remains unclear and several sources of uncertainty in both results have to be addressed.
The presented method for the calculation of hyperfine Stark effects is well suited for future applications on other atomic systems, in particular those with clock transitions involving hole quantum states and/or several unpaired electrons.

### v2

@@ III.4.1 $^{2}F_{7/2}(F=4)$$^{2}F_{5/2}(F=3)$ clock transition
Both the scalar and tensor polarizabilites are numerically invariant to the choice of $M_{J}$ components used for their calculation up to two digits after the decimal point. This is shown for one case using the model QZ/SD33/10au. All scalar and tensor polarizabilites are therefore calculated from $\alpha_{7/2,3/2}$ and $\alpha_{7/2,1/2}$.
For $J=5/2$ states of thulium the inverted equations read
$\alpha_{0}\left({}^{2}F_{5/2}\right)$ $=$ $\frac{4}{3}\alpha_{D}\left({}^{2}F_{5/2,3/2}\right)-\frac{1}{3}\alpha_{D}\left({}^{2}F_{5/2,1/2}\right)$ (27)
$\alpha_{2}\left({}^{2}F_{5/2}\right)$ $=$ $\frac{5}{3}\left[\alpha_{D}\left({}^{2}F_{5/2,3/2}\right)-\alpha_{D}\left({}^{2}F_{5/2,1/2}\right)\right]$ (28)
$\alpha_{0}\left({}^{2}F_{5/2}\right)$ $=$ $\frac{1}{6}\alpha_{D}\left({}^{2}F_{5/2,5/2}\right)+\frac{5}{6}\alpha_{D}\left({}^{2}F_{5/2,3/2}\right)$ (29)
$\alpha_{2}\left({}^{2}F_{5/2}\right)$ $=$ $\frac{5}{6}\left[\alpha_{D}\left({}^{2}F_{5/2,5/2}\right)-\alpha_{D}\left({}^{2}F_{5/2,3/2}\right)\right]$ (30)
This, in turn, allows for calculating the scalar and tensor polarizabilities from the present $M_{J}$-dependent values, the results of which are also given in Table 5.
Table 5: $M_{J}$-dependent static electric dipole polarizabilities $\alpha_{D}$ [a.u.] calculated through the FF method for states $^{M}L_{J,M_{J}}$ where $M=2S+1$ is the spin multiplicity and scalar ($\alpha_{0}$) and tensor ($\alpha_{2}$) polarizabilities for the $^{2}F_{5/2}$ multiplet
$\alpha_{D}$
Model $^{2}F_{5/2,1/2}$ $^{2}F_{5/2,3/2}$ $^{2}F_{5/2,5/2}$ $\alpha_{0}(^{2}F_{5/2})$ $\alpha_{2}(^{2}F_{5/2})$ calc. from $\{M_{J}\}$
TZ/DCHF $186.3443$ $185.0430$ $184.6092$ 2.169-2.169 $1/2,3/2$
TZ/SD15/6au $163.9999$ $162.7315$ $162.3087$ 2.114-2.114 $1/2,3/2$
TZ/SD13_SD15/6au $163.9924$ $162.7252$ $162.3028$ 2.112-2.112 $1/2,3/2$
TZ/SD13_SDT15/6au $161.9244$ $160.8021$ $160.4280$ 1.871-1.871 $1/2,3/2$
TZ/SD13_SDTsppdQ15/6au $167.1381$ $165.8626$ $165.4374$ 2.126-2.126 $1/2,3/2$
QZ/DCHF $186.5082$ $185.1807$ $184.7382$ 2.213-2.213 $1/2,3/2$
QZ/SD15/10au $161.2017$ $159.9065$ $159.4748$ 2.159-2.159 $1/2,3/2$
QZ/SDT15/10au $160.7452$ $159.5866$ $159.2004$ 1.931-1.931 $1/2,3/2$
QZ/SD23/10au $151.1002$ $149.7506$ $149.3007$ 2.249-2.249 $1/2,3/2$
QZ/SD33/10au $151.7513$ $150.3923$ $147.6753$ $149.9393$ 2.265-2.265 $1/2,3/2$
QZ/SD33/10au $149.9395$ 2.264-2.264 $5/2,3/2$
In Ref. NatComm_Tm_2019 Eq. (5) is given
$\Delta\alpha_{\text{DC}}^{s}=\alpha^{s}_{5/2}-\alpha^{s}_{7/2}=-0.063(30)~{\text{a.u.}}$ (31)
from a combination of measurement and calculation. This being a negative quantity means that $\alpha^{s}_{7/2}>\alpha^{s}_{5/2}$.
The electronic ground state $^{2}F_{7/2}$ of the Tm atom can in the Hartree-Fock picture be represented by a $4f^{13}$ configuration written as $4f_{5/2}^{6}\,4f_{7/2}^{7}$ in terms of Hartree-Fock spinors (the $j=5/2$ level is energetically lower than the $j=7/2$ level). In turn, the first excited state $^{2}F_{5/2}$ can be represented as $4f_{5/2}^{5}\,4f_{7/2}^{8}$. According to Ref. desclaux_tables numerical Dirac-Hartree-Fock calculations show that the radial expectation values of the valence spinors are
$\left<\hat{r}\right>_{5/2}$ $=$ $0.763~{\text{a.u.}}$
$\left<\hat{r}\right>_{7/2}$ $=$ $0.780~{\text{a.u.}}$
respectively. This means that, qualitatively, the level with the greater $4f_{7/2}$ occupation is in a straightforward interpretation expected to be the level with the greater static dipole polarizability, since the $4f_{7/2}$ spinors are more diffuse than the $4f_{5/2}$ spinors. The level with the greater $4f_{7/2}$ occupation is the $^{2}F_{5/2}$ level ($8$ electrons occupying the $j=7/2$ spinors). Therefore, in Dirac-Hartree-Fock theory, the expected result is $\alpha^{s}_{5/2}>\alpha^{s}_{7/2}$ which contradicts the result given in Ref. NatComm_Tm_2019 and in Eq. 31 above.
Relativistic many-body calculations yield results for the differential scalar static dipole polarizability defined as $\Delta\alpha_{0}=\alpha_{0}(^{2}F_{5/2})-\alpha_{0}(^{2}F_{7/2})$ and compiled in Table 6.
Table 6: Differential static scalar electric dipole polarizabilities $\alpha_{0}$ [a.u.] for the thulium atom ground term calculated through the FF method from various CI models
Model $\alpha_{0}(^{2}F_{5/2})$ $\alpha_{0}(^{2}F_{7/2})$ $\Delta\alpha_{0}$
TZ/DCHF/0au $184.6092$ $184.5674$ $0.0418$
TZ/SD15/6au $162.3087$ $162.1495$ $0.1592$
TZ/SD13_SD15/6au $162.3028$ $162.1442$ $0.1586$
TZ/SD13_SDT15/6au $160.4280$ $160.3407$ $0.0873$
TZ/SD13_SDTsppdQ15/6au $165.4374$ $165.4144$ $0.0230$
QZ/DCHF/0au $184.7382$ $184.6948$ $0.0434$
QZ/SD15/10au $159.4748$ $159.3583$ $0.1165$
QZ/SDT15/10au $159.2004$ $159.1952$ $0.0052$
QZ/SD23/10au $149.3007$ $149.2431$ $0.0576$
QZ/SD33/10au $149.9393$ $149.8976$ $0.0417$
final $\boldsymbol{-0.134\pm 0.11}$}
Experiment NatComm_Tm_2019 0.063(30)-0.063(30)
It is interesting to note that at all individual levels of calculation the qualitative picture of DCHF theory is reproduced also when electron correlation effects are taken into account. However, when adding the effect of Triple excitations to the model with the greatest number of electrons subjected to the correlation treatment, SD33/10au, the mentioned interpretation at Hartree-Fock level of theory is reverted. Another sizable negative correction from a limited set of quadruple excitations on top of the triple excitations is determined by using the smaller TZ basis set where a calculation of this size becomes feasible. Going beyond this model is not possible with the current implementation due to computational limitations and calculation time.
The final result for the differential static scalar dipole polarizability for the Tm atom $^{2}F_{7/2}(F=4)$$^{2}F_{5/2}(F=3)$ clock transition is obtained by using the result from the model with the greatest number of electrons in the correlation treatment (SD33/10au) as a base value and adding to it corrections due to CI excitation ranks surpassing singles and doubles excitations. In formal terms this calculation reads as
$\Delta\alpha_{0}({\text{final}})$ $=$ $\Delta\alpha_{0}({\text{QZ/SD33/10au}})$
$+\Delta\alpha_{0}({\text{QZ/SDT15/10au}})-\Delta\alpha_{0}({\text{QZ/SD15/10au}})$
$+\Delta\alpha_{0}({\text{TZ/SD13\_SDTsppdQ15/6au}})-\Delta\alpha_{0}({\text{TZ/SD13\_SDT15/6au}})$
Physically speaking, an encompassing treatment of interelectron correlation effects leads to the astonishing result that the static scalar electric dipole polarizability is greater in the $^{2}F_{7/2}$ state than in the $^{2}F_{5/2}$ state of thulium. The simple single-determinant picture for electric polarizability based on Hartree-Fock spinors, therefore, breaks down. The present result confirms qualitatively in this regard the result from Ref. NatComm_Tm_2019 that has been obtained through a combination of experimental measurement and theory. However, due to the sizable corrections found with even the most extensive CI models the uncertainty on the present result must remain rather large. The uncertainty based on the limited treatment of CI excitation ranks, atomic basis sets, correlation effects from inner atomic shells and approximations in the employed atomic Hamiltonian does not compromise the qualitative conclusions.
@@ IV Conclusions
A variational relativistic configuration-interaction approach to the calculation of the hyperfine Stark coefficient in atoms is presented. As a methodological conclusion, the present approach can be applied to electronic transitions of any type in any atom, given that the hyperfine Stark coefficient is calculated as a differential static polarizability.
For the 87Rb atom excitation ranks up to quintuples have been included in the wavefunction expansion and the obtained central value is the theoretical result so far closest to the experimental value. Calculations of similar sophistication for the 133Cs atom yield a final result that is compatible with other high-level theoretical calculations and the 1998 experimental value by Simon et al. Simon_PRA1998 . The present method is then applied to a 169Tm clock transition where so far pure ab-initio calculations have been lacking. The present calculations explain how the sign of the hyperfine Stark coefficient that has previously been measured NatComm_Tm_2019 comes about. The difficulties in obtaining accurate electron correlation effects for relevant properties of 169Tm have also been encountered in the calculation of its ground-state electric quadrupole moment Fleig_PhysRevA.107.032816 .